OpenStudy (gabby1214):

solve using multiplication or division 11= x/9 x=?

OpenStudy (alexandervonhumboldt2):

multiply both sides by 9: x/9*9=11*9 x=11*9=?

OpenStudy (gabby1214):

99?

OpenStudy (mathmale):

That 's a good approach. However, both x/9 and 11 should be within ( ) before you multiply them by 9. This multiplication clears out the fraction.

OpenStudy (mathmale):

Check your own answer by subst. it into the original equation.

OpenStudy (mathmale):

Is 11 = 99/9 true or false? Is your answer correct or not?